OpenStudy (anonymous):

The line segment has endpoints at (3, -8) and (-3, 8). How long is the line segment?

OpenStudy (espex):

Using the Pythagorean theorem you should be able to calculate the distance. From \[c^2 = a^2 + b^2\], where 'a' and 'b' are your (x_0 - x_1) and (y_0 - y_1) respectively, you get the distance formula described as \[d = \sqrt{(x_0 - x_1)^2 + (y_0 - y_1)^2}\] Now simply plug and chug.

OpenStudy (anonymous):

Thank you :)

OpenStudy (espex):

You're welcome.